A 3.05-kg object is moving in a plane, with its x and y coordinates given by x = 8t2 − 4 and y = 5t3 + 4, where x and y are in meters and t is in seconds. Find the magnitude of the net force acting on this object at t = 1.75 s.
Answer #1

Using the Equation v, 16t a, 16 a, -30t 30(1.75 s) 3 2 - 52.5 So acceleration is (52.5) = 54.88m/s2 Net Force is, = ma (3.05k